y''-3y'+2y=3e^-x-10cosx
1
Expert's answer
2020-12-22T20:14:43-0500

"y''-3y'+2y=3 e^x-10 \cos(x)"

"1) y''-3y'+2y=0"

"k^2-3k+2=0"

"D=9-8=1"

"k_1=\frac{3-1}{2}=1"

"k_2=\frac{3+1}{2}=2"

"Y=C_1e^x+C_2e^{2x}"

2) "y^*=(Ax+B)\cdot e^x+C\cdot cos(x)+D\cdot sin(x)"

"(y^*)'=A\cdot e^x+(Ax+B)\cdot e^x-C\cdot sin(x)+D\cdot cos(x)="

"=(Ax+A+B)\cdot e^x-C\cdot sin(x)+D\cdot cos(x)"

"(y^*)''=(Ax+2A+B)\cdot e^x-C\cos(x)-D\cdot \sin(x). Then:"

"(Ax+2A+B)\cdot e^x-C\cos(x)-D\cdot \sin(x)-3(Ax+A+B)\cdot e^x+3C\cdot sin(x)-3D\cdot cos(x)+2(Ax+B)\cdot e^x+2C\cdot cos(x)+2D\cdot sin(x)=3e^x-10cos(x)"

"(Ax+2A+B-3Ax-3A-3B+2Ax+2B)e^x+\cos(x) (-C-3D+2C)+\sin(x)(-D+3C+2D)=3e^x-10cos(x)"

"A=-3, B=const"

"C-3D=-10, 3C+D=0 ~then ~D=3, C=-1;"

"y^*=(B-3x)e^x-\cos(x)+3\sin(x)"

"y=Y+y^*=C_1e^x+C_2e^{2x}+(B-3x)e^x-\cos(x)+3\sin(x)="

"=e^x(C^*-3x)+C_2e^{2x}-\cos(x)+3\sin(x), C^*-const"